# Investigating Experiential Effects in Online Chess using a Hierarchical Bayesian Analysis

### v1

@@ 1 Introduction
We find little evidence for strong persistent winner-loser effects, with this conclusion consistent across the range of players and abilities analysed. Individuals appear to show some variability, and there is some evidence that some players demonstrate such effects. However, these effects do not appear to be systematic.
This article is organized as follows. We first provide a brief history of the study of these effects across disciplines in Section 1.1. In Section 2 we describe in detail the data we use here, highlighting key properties which make it well suited for the study of winner-loser effects. Section 3 gives the proposed hierarchical Bayesian regression model, detailing the model and prior specification. In Section 4 we fit the proposed model, highlighting the results for two cohorts of chess players of differing abilities. Section 5 provides a detailed discussion of several procedures used to ensure our model is able to correctly capture the temporal structure present in such data. Finally, in Section 6 we summarise the results of this model and highlight important considerations for future work.
@@ 1.1 A history of the study of experiential effects in human and animal competition
Notably, there has been a relative lack of investigation of human competitive interactions that incorporate thousands of competitive interactions over a longer period of time such as is available with data from online chess matches. Similarly, the existing models for examining such effects are often limited or somewhat domain specific. In this work we develop a natural statistical model for such interaction data which can be applied to any dataset of repeated competition. As discussed below, it can also incorporate information about these competitions as available.
@@ 2 Data
In this work we use rated standard games played on Lichess, a popular open source chess server. Lichess and other chess platforms pair random opponents who are believed to have a similar chess ability, to create games where both players have an opportunity to win. In online and over the board chess, it is common to estimate the ability of an individual using a numeric rating based on their past performance. When an individual creates an account on Lichess, they are a given a default rating of 1500 across all chess formats. As the player plays rated games at a specific time format, their rating will update based on their performance and the rating of their opponents. Lichess uses the Glicko-2 rating system [24], which also estimates a measure of variability of this rating. This variability decreases as more games are played and the estimated rating becomes a more accurate measure of a players ability at that time format [23]. In Figure 1 we show the rating evolution of two players with a similar number of games played. We see that in each case their rating improves over time, with significant variability before somewhat stabilizing.
@@
Lichess allows a wide range of variants and time formats of chess, with games ranging from several minutes to several days. The only variant we consider are standard chess formats, which is the vast majoirty of games on Lichess. In these time controlled chess games, each player starts with the same amount of time to make their moves. The two most common formats are ``Bullet'' (games where each player has 1-2 minutes) and ``Blitz'' (games where each player has 3-5 minutes). These short games commonly end with one player winning via checkmate, resignation or running out of time. Draws are possible but relatively rare in these formats. We consider only the most played time control in bullet and blitz for analysis: 60+0 (60 seconds for all moves with no increment) and 180+0 (3 minutes for all moves with no increment) respectively.
@@
We wish to examine the presence of experiential effects across a range of chess abilities and playing history. Using a players current rating, we collected data for 40 players in 3 seperate rating cohorts each: 1700-1900, 2000-2200 and 2300-2500. Between October 23-31, 2023, we randomly selected 40 players within each of these rating ranges that had recently played bullet games of any format, and scraped information about the games they had played up to that date. These rating cohorts correspond to a range of moderate to high chess ability [^2] Lichess provides a weekly distribution of ratings on the platform. At the time of writing, 47% of bullet players had a rating lower than 1500, while only 1.5% had a rating higher than 2500.. We also collect all games played by 25 grandmaster (GM) accounts on Lichess in early 2024. Grandmaster is an official title awarded to players who demonstrate excellent ability in chess. At the time of writing, this extremely selective title has only been awarded to 2087 players in history. For the GMs we selected, the average rating was 3055 for Bullet and 2854 for Blitz games. We denote the 40 players in each rating cohort and the 25 GMs as ``focal'' players in this paper.
@@
For every game by each focal player in the 60+0 and 180+0 time controls, we observe their current rating, the rating of their opponent, the colour of each player and the outcome of the game. For brevity, we focus our findings on the 1700-1900 and GM cohorts in what follows. The results obtained for the other cohorts, unless noted, agree with those presented below, with some of the corresponding analyses included in the supplementary materials.
@@ 2.1 Data Summary
For the 1700-1900 cohort, 36 focal players played over 590,000 60+0 Rated Bullet Games, more than 87% of all rated bullet games they play, with each playing on average over 16000 games. For the games played by grandmasters, 75% of all bullet games are 60+0, corresponding to more than 220,000 games, with each of the 25 players playing an average of almost 9000 games. For blitz games, we have 107,000 180+0 Rated Blitz Games in the 1700-1900 cohort (58% of all blitz games), with 25 focal players playing an average of 4140 games. For the GM cohort we have 33,000 games (98% of all blitz games), with each playing on average 1292 games.
Fig. 1: Evolution of Lichess Bullet ratings for one player in the 1700-1900 cohort and another in the GM cohort.
@@
Fig. 2: Time gaps between successive bullet games (left) and the distribution of the lengths of a session of bullet games (right) for the GM cohort.
In Figure 2 we show the time between the start of successive bullet games in the GM cohort on a log scale. The maximum length of these games is 2 minutes. Across all grandmaster games, 88.6% of the games are played within 5 minutes of the previous game. This indicates that periods where players play more than one game in a row are very common (for the 1700-1900 cohort this is 85% of all bullet games).
Based on this information we define bullet games played in the same session as those which start within 5 minutes of the start of the previous game. With this definition we can then determine how many games are played in a given session. We plot the distribution of the number of games in a session, from the GM cohort, in Figure 2. The median number of games played in such a session is 5, with 78% of all sessions involving more than 1 game. It is very common for an individual to play multiple games in quick succession. It is commonly thought that experiential effects are short term, with the result of the previous competition having an impact on the next [17]. If experiential effects do occur, it seems plausible that they would be detected in games of this form, where players who win or lose can almost immediately enter another contest.
While it is believed that winner-loser effects occur based on the most recent history, we wish to investigate if longer term effects can exist also. In the model described below we will investigate the presence of these effects based on the result immediately proceeding the current game only. However, we will also investigate if these effects have a longer duration by using the performance in the previous 10 games as a predictor of the current outcome. It is still relatively common for players to play 10 or more games in succession, with 26% of sessions played by GMs, as we have defined them, containing more than 10 games [^3] We see similar proportions for the 1700-1900 cohort.. This will therefore often include periods where players have a gap between successive games. Fitting our proposed model taking account of this longer history allows investigation of more prolonged experiential effects.
@@ 3 Modeling Winner-Loser Effects with a Hierarchical Logistic Regression
Here $p_{ij}$ is the probability of the $j$-th focal player winning their $i$-th game, $y_{ij}$. $\alpha_{j}$ is a player level random effect, indicating a players individual ability to win a game against an equally rated opponent when playing with the black pieces. $\beta_{j}$ is a player level random effect, accounting for the win ratio $\tilde{x}_{ij}^{n}$ of the focal player over their previous $n$ games. We discuss this in more detail below. $\bm{\gamma}$ is a vector of fixed effects corresponding to two individual game covariates, the colour played by the focal player and the difference in rating between the focal player and their opponent. We partially pool the $\alpha$ and $\beta$ coefficients with a bivariate normal prior distribution and place independent normal priors on the fixed effects. The prior specification for all parameters is given by
@@
$\Omega\sim\mbox{LKJcorr}(2).$
We normalize the win ratio by the average win ratio of player $j$ across all games they play in the dataset. If $x_{ij}^{n}$ is the proportion of games that the $j$-th player has won in the past $n$ games then $\tilde{x}_{ij}^{n}=x_{ij}^{n}-\bar{x}_{j}$, where $\bar{x}_{j}$ is their overall win proportion for games of that format. $\beta_{j}$ captures how a players win probability changes as their recent performance deviates from their overall average win ratio.
@@
We decompose the covariance matrix $\Sigma$ into a correlation matrix $\Omega$ and the individual scale of the two random effects. For the correlation matrix we choose an LKJ prior with shape parameter 2 [32]. This is a prior over the set of correlation matrices with the chosen shape parameter favoring less correlation. We use independent Inverse-Gamma priors with shape and scale 1 for all scale parameters. In initial model fitting we found that less informative priors for the scale parameters gave equivalent results while making posterior convergence more challenging.
This model is a natural approach to investigating this phenomenon in online chess, and is similar to models which have been considered in both the sport analytics and animal behaviour literature [11, 17]. The flexibility of this model allows it to be applied to a range of potential datasets consisting of repeated competition observed for a number of individuals.
@@ 4 The Fitted Model
We fit the proposed Bayesian model with Stan [45], using the CmdStanR [18] interface in R [41]. We use default settings for fitting these models, running 4 chains each with 1000 warmup and 1000 sampling iterations. We fit the proposed model for each of the four cohorts discussed previously. In each case the models indicate posterior convergence, as indicated by $\hat{R}$ values close to 1 and large effective sample sizes [22]. We discuss the results of these models for two cohorts: the 1700-1900 cohort and the GM cohort. Similar results for the other cohorts are included in the supplementary material. All code required to reproduce this analysis is available in the associated Github repository.
@@
We first fit the proposed model for standard bullet games as described above with $n=1$, utilising the previous game as the only history. We first examine the key global parameters ($\gamma_{1},\gamma_{2},\mu_{\beta}$), in all four cohorts, with posterior credible intervals for each shown in Figure 3. The main difference is seen in the fixed effects $\gamma_{1}$ and $\gamma_{2}$. The estimated posterior mean of the colour effect ($\gamma_{1}$) is 0.091 and 0.25 for the 1700-1900 and GM cohorts respectively. For a player with a win probability of 0.5 playing as black against an equally rated opponent, playing as white would increase the win probability on average to 0.56 for the GMs but only 0.52 for the 1700-1900s. Similarly, the estimated posterior mean of the rating effect ($\gamma_{2}$) is 0.0037 and 0.0053 for the 1700-1900 and GM cohorts respectively. For a player with a win probability of 0.5 playing as black against an equally rated opponent, playing against an opponent rated 100 points lower increases the win probability on average to 0.63 for GMs and 0.59 for 1700-1900s. This agrees with the common beliefs that such effects have a larger impact in higher rated players [^4] Multiple references are given in [49] and that colour assignment and differences in skill play a larger role in determining the winner of a game at the professional level. In both cohorts the posterior mean value for $\mu_{\beta}$ is centered around 0, indicating that there is little evidence for a global experiential effect. We also include the estimates for each of these parameters for the two intermediate cohorts in Figure 3. We see a clear linear pattern in the estimates of $\gamma_{1}$ and $\gamma_{2}$, corresponding to increasing ability in the cohort. However, for all cohorts, the estimate of $\mu_{\beta}$ clearly centered around 0.
Fig. 3: The 66% and 95% credible intervals of the estimated colour ($\gamma_{1}$), rating ($\gamma_{2}$) and global experiential ($\mu_{\beta}$) effects in the bullet games. The posterior means of the colour and rating effects increase with skill. There is little evidence for a global winner effect in any cohort.
@@
We show histograms and credible intervals of the posterior estimates of the change in win probability for 36 focal players in the 1700-1900 cohort [^5] Four players in the 1700-1900 cohort were excluded due to not playing any 60+0 bullet games. and 25 in the GM cohort. We see that many of these players have changes close to 0. The posterior means for this quantity across these two cohorts range from -0.10 to 0.12. The majority of 95% posterior intervals contain 0 (or being close to 0). These effects are small, and indicate small changes in the win probability based on performance in the previous game. One interesting observation here is that the player with the smallest estimated change in the GM cohort, the username DrNykerstein, is one of the accounts of former world champion Magnus Carlsen. While this result indicates that Carlsen is less likely to win the next game if he had won the previous game than other players, there are several potential factors which could explain this. One possible explanation concerns data quality. Carlsen is widely known for playing in an ``erratic" manner in casual games online, and as seen in live streams, may not be taking such games seriously. In terms of the model used, it is possible that Carlsen is different to other players and (for example) has a winner and loser effect of different sizes. This would not be captured by our current model.
Fig. 4: Interpretation of the estimated player-level experiential effects ($\beta_{j}$ for player $j$). The change in win probability represents the additive boost in win probability a player receives from the experiential effect when changing their previous game from a loss to a win. The 66% and 95% credible intervals are shown. Most players exhibit little evidence for non-zero experiential effects, which results in changes in the win probability that are close to zero. There is a slightly larger proportion of GMs showing evidence of possessing such effects compared to the 1700-1900 cohort.
@@
Our proposed model shows little evidence for winner-loser effects above when $n=1$, which corresponds to the result of the previous game alone impacting subsequent performance. However, it is possible that experiential effects could require longer time periods to develop. To investigate this we refit the above model now using $n=10$ for computing the win ratio $\tilde{x}_{ij}^{n}$. It is relatively common for sessions involving more than 10 games to be played consecutively, with more than 20% of all sessions across both cohorts containing at least 11 games. In Figure 5(a) we show the posterior estimates for the three key model parameters for both $n=1$ and $n=10$ for the GM cohort. We note that the posterior distributions for all parameters except for $\tau_{2}$ are essentially identical. In both cases our posterior distribution for $\mu_{\beta}$ is centered around 0, indicating there is no evidence for a global effect. The posterior distribution of $\tau_{2}$ is shifted towards larger values due to the increased variability of $\tilde{x}_{ij}^{n}$ when $n=10$. Individual winner-loser effects (shown in the supplementary material) show similar patterns with increased variability when $n=10$, with a similar proportion of posterior intervals containing 0 in both cases. It does not appear that a longer term history provides additional evidence for the presence of strong or persistent experiential effects.
@@
The previous analysis has investigated the presence of experiential effects in short bullet games lasting at most two minutes. To consider the possibility that longer competition is required for such effects to form, we repeat the previous analysis now using blitz games. These are games of up to 6 minutes in length. Although these games are longer, it is still common for successive sequences of games. We define a focal player starting another blitz game within 7 minutes of their previous game as games played in the same session. With this definition, two thirds of sessions played in the GM cohort contain more than one game, with almost three quarters of games in the 1700-1900 cohort. As described in Section 2, we have substantially more bullet than blitz games for all players.
We compare the fit of our proposed model with $n=1$ for the GM cohort. The posterior estimates of key parameters are shown in Figure 5(b). The estimated fixed effects $\gamma_{1},\gamma_{2}$ show larger variability in the smaller blitz dataset. In both cases, we see the distribution of the estimated global winner-/loser-effect is centered at 0 with the same variability for both datasets.
@@
Fig. 5: Comparison of the estimated fixed and global effects in varying settings for the GM cohort. In each case we show 66% and 95% posterior intervals.
@@
Fig. 6: Winner-loser effects for the GM data, from fitting the proposed model to Bullet and Blitz data. We see considerably larger variability with the smaller blitz dataset.
Overall, we see little evidence for consistent experiential effects from fitting the proposed Bayesian logistic regression model to large volumes of online chess data. Across a variety of formats and time scales for the formation of these effects, we do not see any evidence for such effects across the population of players considered. While there is some evidence for individual players displaying winner-loser effects, these appear to be weak. An important concern with the current analysis, investigating these effects in the context of recent performance, is the potential for high temporal correlation between successive games. In Section 5 we consider this and provide several tools to validate our model.

### v2

@@ 1 Introduction
We find little evidence for strong persistent winner-loser effects, with this conclusion consistent across the range of players and abilities analysed. Individuals appear to show some variability, and there is some evidence that some players demonstrate such effects. While there is somewhat increased evidence for these winner-loser effects in the top international competitors, these effects do not appear to be systematic across the population of chess players analysed.
This article is organized as follows. We first provide a brief history of the study of these effects across disciplines in Section 1.1. In Section 2 we describe in detail the data we use here, highlighting key properties which make it well suited for the study of winner-loser effects. Section 3 gives the proposed hierarchical Bayesian regression model, detailing the model and prior specification. In Section 4 we fit the proposed model, highlighting the results for two cohorts of chess players of differing abilities. Section 5 provides a detailed discussion of one natural procedure used to ensure our model is able to correctly capture the temporal structure present in such data. We include several further model checking approaches in the appendix. Finally, in Section 6 we summarise the results of this model and highlight important considerations for future work.
@@ 1.1 A history of the study of experiential effects in human and animal competition
Notably, there has been a relative lack of investigation of human competitive interactions that incorporate thousands of competitive interactions over a longer period of time such as is available with data from online chess matches. Similarly, the existing models for examining such effects are often limited or somewhat domain specific. While models such as [34] examine experiential effects in animals, their approach is largely exploratory, examining the relationship between several potential outcomes of interest and previous performance using correlation and generalized linear mixed models. This form of approach, along with traditional statistical tests, is common in the animal literature. In sports, models which have been proposed cannot easily be modified to other scenarios. While an advanced statistical model was proposed to investigate the hot hand in professional darts [41], this utilises the fact that each player has three throws in each turn, examining experiential effects both within a turn and across a leg (several turns). It is unclear how to extend this to more general forms of competition or interactions. Similarly, [20] use causal inference techniques to examine the impact of teams scoring before half time in soccer, matching these goals to similar attempts which instead hit the post. However, it may be challenging to modify this matching approach to other games or scenarios. In this work we develop a natural statistical model for such interaction data which can be applied to any dataset of repeated competition. As discussed below, it can also incorporate information about these competitions as available.
@@ 2 Data
In this work we use rated standard games played on Lichess, a popular open source chess server. Lichess and other chess platforms pair random opponents who are believed to have a similar chess ability, to create games where both players have an opportunity to win. In online and over the board chess, it is common to estimate the ability of an individual using a numeric rating based on their past performance. When an individual creates an account on Lichess, they are a given a default rating of 1500 across all chess formats. As the player plays rated games at a specific time format, their rating will update based on their performance and the rating of their opponents. Lichess uses the Glicko-2 rating system [24, 26], which can also estimate a measure of variability of this rating. This new system builds on the original system of [23] by producing larger variability in the estimated ratings if unexpected results occur. In Figure 1 we show the rating evolution of six players with a similar number of games played from the two cohorts we analyse in detail below. We include three players who are among the highest rated on Lichess (two of whom are known elite players). We also include three players of a more moderate ability. These players show some of the typical behaviour across the range of players we consider. These ratings can show large changes initially, with smaller changes present after several hundred games are played. Similarly, some of these players, such as ``SeanBambic", can show a gradual upward trend, indicating gradual improvement in skill over time.
Figure 1: Evolution of Bullet Glicko-2 ratings for 3 players in the GM cohort and 3 players in the 1700-1900 cohort. ``DrNykterstein" (Magnus Carlsen) and ``alireza2003" (Alireza Firouzja) are among the current top players in the world, with Carlsen being at the number 1 spot since 2011.
@@
Lichess allows a wide range of variants and time formats of chess, with games ranging from several minutes to several days. The only variants we consider are standard chess formats, which is the vast majority of games on Lichess. In these time controlled chess games, each player starts with the same amount of time to make their moves. The two most common formats are ``Bullet'' (games where each player has 1-2 minutes) and ``Blitz'' (games where each player has 3-5 minutes). These short games commonly end with one player winning via checkmate, resignation or running out of time. Draws are possible but relatively rare in these formats. We consider only the most played time controls in bullet and blitz for analysis: 60+0 and 180+0. The first number in these time controls represents how many seconds each player starts with to make all their moves. The second represents the increment, which is the number of additional seconds added to their remaining time after making each move.
@@
We wish to examine the presence of experiential effects across a range of chess abilities and playing history. Using a players current rating, we collected data for 40 players in each of 3 separate rating cohorts: 1700-1900, 2000-2200 and 2300-2500. Between October 23-31, 2023, we randomly selected 40 players within each of these rating ranges that had recently played bullet games of any format, and scraped information about the games they had played up to that date. These rating cohorts correspond to a range of moderate to high chess ability [^2] Lichess provides a weekly distribution of ratings on the platform. At the time of writing, 47% of bullet players had a rating lower than 1500, while only 1.5% had a rating higher than 2500.. We also collect all games played by 25 grandmaster (GM) accounts on Lichess in early 2024. Grandmaster is an official title awarded to players who demonstrate excellent ability in chess. At the time of writing, this extremely selective title has only been awarded to 2087 players in history. For the GMs we selected, the average rating was 3055 for Bullet and 2854 for Blitz games. We denote the 40 players in each rating cohort and the 25 GMs as ``focal'' players in this paper. We have made this data, along with the code used to scrape it, publicly available in our associated Github repository.
@@
For every game by each focal player in the 60+0 and 180+0 time controls, we observe their current rating, the rating of their opponent, the colour of each player and the outcome of the game. Importantly, we note that we are unable to obtain the uncertainty in a player's rating (known as the rating deviation), which is a key component in the Glicko-2 rating system. For brevity, we focus our findings on the 1700-1900 and GM cohorts in what follows. The results obtained for the other cohorts, unless noted, agree with those presented below, with some of the corresponding analyses included in the appendix.
@@ 2.1 Data Summary
We summarize the total number of games we use for both the 1700-1900 and GM cohort in Table 1. For the 1700-1900 cohort we have 36 focal players playing 60+0 Bullet Games. This is more than 87% of all rated bullet games they play (across all possible time limits and increments), with each averaging over 16000 games. For the grandmasters 60+0 games correspond to 75% of all bullet games played, averaging over 8000 per player. The corresponding statistics for 180+0 Blitz games are shown in Table 1. Bullet games are considerably more popular across all cohorts, perhaps due to the short time required.
Game Type Cohort Num. Players Games (count) Percent of Games Avg Games per Player
Bullet (60+0) 1700–1900 36 590,253 87% 16,396
Bullet (60+0) Grandmasters 23 196,522 75% 8,544
Blitz (180+0) 1700–1900 36 107,670 58% 4,141
Blitz (180+0) Grandmasters 23 34,251 98% 1,489
Table 1: Summary statistics for bullet and blitz games by cohort and format. Here percent of Games corresponds to how many of all Bullet/Blitz games were of this specific time format.
@@
Figure 2: Time gaps between successive bullet games (left) and the distribution of the lengths of a session of bullet games (right) for the GM cohort.
In Figure 2(left) we show the time between the start of successive bullet games in the GM cohort on a log scale. The maximum length of these games is 2 minutes. Across all grandmaster games, 88.6% of the games are played within 5 minutes of the previous game. This indicates that periods where players play more than one game in a row are very common (for the 1700-1900 cohort this is 85% of all bullet games).
Based on this information we define bullet games played in the same session as those which start within 5 minutes of the start of the previous game. With this definition we can then determine how many games are played in a given session. We plot the distribution of the number of games in a session, from the GM cohort, in Figure 2(right). The median number of games played in such a session is 4, with 76.3% of all sessions involving more than 1 game. It is very common for an individual to play multiple games in quick succession. It is commonly thought that experiential effects are short term, with the result of the previous competition having an impact on the next [17]. If experiential effects do occur, it seems plausible that they would be detected in games of this form, where players who win or lose can almost immediately enter another contest.
While it is believed that winner-loser effects occur based on the most recent history, we wish to investigate if longer term effects can exist also. In the model described below we will investigate the presence of these effects based on the result immediately proceeding the current game only. However, we will also investigate if these effects have a longer duration by using the performance in the previous 10 games in a session as a predictor of the current outcome. It is still relatively common for players to play 10 or more games in succession, with 26.3% of sessions played by GMs, as we have defined them, containing more than 10 games [^3] We see similar proportions for the 1700-1900 cohort.. Fitting our proposed model taking account of this longer history allows investigation of more prolonged experiential effects.
@@ 3 Modeling Winner-Loser Effects with a Hierarchical Logistic Regression
Here $p_{ij}$ is the probability of the $j$-th focal player winning their $i$-th game, $y_{ij}$. $\alpha_{j}$ is a player level random effect, indicating a player's individual ability to win a game against an equally rated opponent when playing with the black pieces. $\beta_{j}$ is a player level random effect, accounting for the win ratio $\tilde{x}_{ij}^{n}$ of the focal player over their previous $n$ games. We discuss this in more detail below. $\bm{\gamma}$ is a vector of fixed effects corresponding to two individual game covariates, the colour played by the focal player and the difference in rating between the focal player and their opponent. We partially pool the $\alpha$ and $\beta$ coefficients with a bivariate normal prior distribution and place independent normal priors on the fixed effects. The prior specification for all parameters is given by
@@
$\Omega\sim\mbox{LKJcorr}(2),$
where $\mbox{LKJcorr}(2)$ is a Lewandowski-Kurowicka-Joe (LKJ) distribution with shape parameter 2 [35] and $\mathcal{N}_{+}(0,1)$ is a positively truncated standard normal distribution.
We normalize the win ratio by the cumulative win ratio of player $j$ up to the the $i$-th game. If $x_{ij}^{n}$ is the proportion of games the $j$-th player has won in the past $n$ games then $\tilde{x}_{ij}^{n}=x_{ij}^{n}-\bar{x}_{j}$, where $\bar{x}_{j}$ is their cumulative win proportion up to the current game. Importantly, we consider previous games to only be those played in the current session, as defined previously. If this is the first game in a session then we define $\tilde{x}_{ij}^{n}=0$. Note that if $n=1$, meaning we only consider the previous game in the session, then $\tilde{x}_{ij}^{n}$ will be positive if the previous game was a victory, and negative if the previous game was a loss.
@@
We decompose the covariance matrix $\Sigma$ into a correlation matrix $\Omega$ and the individual scale of the two random effects. For the correlation matrix we choose a Lewandowski-Kurowicka-Joe (LKJ) distribution as the prior, with shape parameter 2 [35]. This is a prior over the set of correlation matrices with the chosen shape parameter favoring less correlation. We use independent half standard normal priors for the scale parameters $\tau_{1},\tau_{2}$. In initial model fitting we found that less informative priors for the scale parameters, such as Inverse-Gamma, gave equivalent results while making posterior convergence more challenging.
This model is a natural approach to investigating this phenomenon in online chess, and builds on models which have been considered in both the sport analytics and animal behaviour literature. [17] examine experiential effects in primates by modifying the ELO rating system. They modify the K-factor in this model, estimating this as a function of contest level predictors. This model could be applied to similar scenarios to the one we consider, however the ELO rating has largely been superseded by the Glicko-2 method. Our Bayesian approach also provides additional uncertainty quantification for identifying experiential effects. The model for hockey goalie performance of [11] is similar to our approach, in that they fit a hierarchical logistic regression to contest outcomes and include historic performance as a predictor. Our model expands on this approach by considered a more flexible modeling structure, specifically designed to identify a global experiential effect. We also develop novel model checking methods. The flexibility of our model also allows it to be applied to a range of potential datasets consisting of repeated competition observed for a number of individuals.
@@ 4 The Fitted Model
We fit the proposed Bayesian model with Stan [48], using the CmdStanR [18] interface in R [44]. We use default settings for fitting these models, running 4 chains each with 1000 warmup and 1000 sampling iterations. We fit the proposed model for each of the four cohorts discussed previously. In each case the models indicate posterior convergence, as indicated by $\hat{R}$ values close to 1 and large effective sample sizes [22]. We discuss the results of these models for two cohorts: the 1700-1900 cohort and the GM cohort. Similar results for the other cohorts are included in the appendix. All code required to reproduce this analysis will be available in the associated Github repository, along with access to the data used.
@@
We first fit the proposed model for standard bullet games as described above with $n=1$, utilising the previous game in the same session as the only history. We examine the key global parameters ($\gamma_{1},\gamma_{2},\mu_{\beta}$), in all four cohorts, with posterior credible intervals for each shown in Figure 3. The main difference is seen in the fixed effects $\gamma_{1}$ and $\gamma_{2}$. The estimated posterior mean of the colour effect ($\gamma_{1}$) is 0.087 and 0.23 for the 1700-1900 and GM cohorts respectively. For a player with a win probability of 0.5 playing as black against an equally rated opponent, playing as white would increase the win probability on average to 0.56 for the GMs but only 0.52 for the 1700-1900s. This agrees with common beliefs for the first move advantage for GMs. Among decisive results in top GM games, white wins 54-56% of games, compared to black winning 44-46% of games. It is also commonly believed that this advantage diminishes at lower levels of play [29]. Similarly, the estimated posterior mean of the rating effect ($\gamma_{2}$) is 0.0038 and 0.0053 for the 1700-1900 and GM cohorts respectively. For a player with a win probability of 0.5 playing as black against an equally rated opponent, playing against an opponent rated 100 points lower increases the win probability on average to 0.63 for GMs and 0.59 for 1700-1900s. The increased impact of these effects in higher rated players seems reasonable, with elite players better able to capitalize on these advantages.
In the 1700-1900 cohort the posterior mean value for $\mu_{\beta}$ is centered around 0, indicating that there is little evidence for a global experiential effect. The posterior mean estimate for the GM cohort is positive, although the 95% interval contains 0, indicating limited weak evidence for a positive experiential effects. We also include the estimates for each of these parameters for the two intermediate cohorts in Figure 3. We see a clear linear pattern in the estimates of $\gamma_{1}$ and $\gamma_{2}$, corresponding to increasing ability in the cohort. However, for all cohorts, the estimate of $\mu_{\beta}$ is centered close to 0.
Figure 3: The 66% and 95% credible intervals of the estimated colour ($\gamma_{1}$), rating ($\gamma_{2}$) and global experiential ($\mu_{\beta}$) effects in the bullet games. The posterior means of the colour and rating effects increase with skill. There is little evidence for a global winner effect in any cohort.
@@
We show credible intervals of the posterior estimates of the change in win probability for 36 focal players in the 1700-1900 cohort [^4] Four players in the 1700-1900 cohort were excluded due to not playing any 60+0 bullet games. and 25 in the GM cohort. We see that many of these players have changes close to 0. The posterior means for this quantity across these two cohorts range from -0.02 to 0.03. The majority of 95% posterior intervals contain 0 (or are close to 0). These effects are small, and indicate small changes in the win probability based on performance in the previous game. One interesting observation here is that the player with the second smallest estimated change in the GM cohort, the username DrNykerstein, is one of the accounts of former world champion Magnus Carlsen. While this result seems to indicate that Carlsen is less likely to win the next game if he had won the previous game than other players, there are several potential factors which could explain this. One possible explanation concerns data quality. Carlsen is widely known for playing in an ``erratic" manner in casual games online, and as seen in live streams, may not be taking such games seriously. In terms of the model used, it is possible that Carlsen is different to other players and (for example) has a winner and loser effect of different sizes. This would not be captured by our current model.
Figure 4: Interpretation of the estimated player-level experiential effects ($\beta_{j}$ for player $j$) for the 1700-1900 Cohort (left) and GM cohort (right). The change in win probability represents the additive boost in win probability a player receives from the experiential effect when changing their previous game from a loss to a win. The 66% and 95% credible intervals are shown. Most players exhibit little evidence for non-zero experiential effects, which results in changes in the win probability that are close to zero.
@@
Our proposed model shows little evidence for winner-loser effects above when $n=1$, which corresponds to the result of the previous game alone impacting subsequent performance. However, it is possible that experiential effects could require longer time periods to develop. To investigate this we refit the above model now using $n=10$ for computing the win ratio $\tilde{x}_{ij}^{n}$. It is relatively common for sessions involving more than 10 games to be played consecutively, with more than 20% of all sessions across both cohorts containing at least 11 games. In Figure 5(a) we show the posterior estimates for the three key model parameters for both $n=1$ and $n=10$ for the GM cohort. We note that the posterior distributions for all parameters except for $\tau_{2}$ are essentially identical with slightly more uncertainty with $n=10$, due to the decreased number of sessions. The posterior mean estimate for $\mu_{\beta}$ is closer to 0 when $n=10$, although the intervals show substantial overlap. The posterior distribution of $\tau_{2}$ is shifted towards larger values due to the increased variability of $\tilde{x}_{ij}^{n}$ when $n=10$. Individual winner-loser effects (shown in the supplementary material) show similar patterns with increased variability when $n=10$, with a similar proportion of posterior intervals containing 0 in both cases. It does not appear that a longer term history provides additional evidence for the presence of strong or persistent experiential effects.
@@
The previous analysis has investigated the presence of experiential effects in short bullet games lasting at most two minutes. To consider the possibility that longer competition is required for such effects to form, we repeat the previous analysis now using blitz games. These are games of up to 6 minutes in length. Although these games are longer, it is still common for successive sequences of games. We define a focal player starting another blitz game within 7 minutes of their previous game as games played in the same session. With this definition, two thirds of sessions played in the GM cohort contain more than one game, along with almost three quarters of games in the 1700-1900 cohort. As described in Section 2, we have substantially more bullet than blitz games for all players.
We compare the fit of our proposed model with $n=1$ for the GM cohort. The posterior estimates of key parameters are shown in Figure 5(b). The estimated fixed effects $\gamma_{1},\gamma_{2}$ show larger variability in the smaller blitz dataset. In both cases, we see the distribution of the estimated global winner-/loser-effect is close to zero, with significantly more variability in the smaller Blitz dataset.
@@
Figure 5: Comparison of the estimated fixed and global effects in varying settings for the GM cohort. In each case we show 66% and 95% posterior intervals.
@@
Figure 6: Winner-loser effects for the GM data, from fitting the proposed model to Bullet and Blitz data. We see considerably larger variability with the smaller blitz dataset.
Overall, we see little evidence for consistent experiential effects from fitting the proposed Bayesian logistic regression model to large volumes of online chess data. Across a variety of formats and time scales for the formation of these effects, we do not see any evidence for such effects across the population of players considered. While there is some evidence for individual players displaying winner-loser effects, and some evidence of a small global effect in the highest rated players, these effects appear to be weak. An important concern with the current analysis, investigating these effects in the context of recent performance, is the potential for high temporal correlation between successive games. In Section 5 we consider this and provide several tools to validate our model.